# Mod $\ell$ non-vanishing of self-dual Hecke $L$-values over CM fields and applications

### v1

@@ 1.4. Strategy
The non-vanishing problems (Q1) and (Q2) concerns the arithmetic of a self-dual Hecke character or that of $U(1)$ over $K$. Via theta correspondence, we recast them as problems on a totally definite unitary group $U(2)$. The latter turn out to be amenable to automorphic tools.
@@ 1.4.1. Shimura set
For a CM field $K$ and $F$ the maximal totally real subfield, we have a naturally associated totally definite quaternion algebra $B$ over $F$ characterised by
@@
for any place $v$ of $F$. Note that $K$ embeds into $B$ as an $F$-algebra, and we fix such an apt embedding.
@@ 1.4.2. Ancillary results
Our mod $\ell$ non-vanishing of Hecke $L$-values is based on Theorems 1.11 and 1.12 below, which concern the arithmetic of the Shimura set.
For a self-dual Hecke character $\lambda$ over $K$ of infinity type $\Sigma+\kappa(1-c)$, let $\phi_{\lambda}$ be the associated ${\mathrm{GL}}_{2}(F)$-theta series of weight $k:=2\Sigma+2\kappa$ and level $U_{0}(\mathfrak{N}_{\lambda})$ for $\mathfrak{N}_{\lambda}:=D_{K/F}{\rm N}_{K/F}({\mathrm{Cond}}(\lambda))$. Write $\mathfrak{N}_{\lambda}=\mathfrak{N}_{\lambda}^{+}\mathfrak{N}_{\lambda}^{-}$ with $\mathfrak{N}_{\lambda}^{+}$ (resp. $\mathfrak{N}_{\lambda}^{-}$) only divisible by split (resp. non-split) primes in $K$. Let $\pi_{\lambda}$ be the Jacquet–Langlands transfer of the automorphic representation generated by $\phi_{\lambda}$ to $B^{\times}$. Let $\ell$ be a prime such that $(\ell,6{\mathfrak{p}}D_{K})=1$. Let
@@
be the CM form as in subsection 3.4. It is a test vector in the sense of Gross and Prasad [18], which is $\ell$-primitive and $K_{v}^{\times}$-invariant for all places $v|\mathfrak{N}_{\lambda}^{-}$. We emphasise that $f_{\lambda}$ is not a newform [^5] In this setting the finite part of the discriminant of $B$ divides $D_{K/F}$. Consequently, at places dividing $D_{K/F}$, newform is not a test vector under an optimal embedding $K\hookrightarrow B$..
@@
The Rankin–Selberg $L$-values maybe studied via the Waldspurger formula. While employing it, the period
@@
arises naturally, where $(\phi_{\lambda},\phi_{\lambda})_{U_{0}(\mathfrak{N}_{\lambda})}$ and $\langle{f_{\lambda},f_{\lambda}}\rangle_{\widehat{R}^{\times}}$ are Petersson norms. In view of the explicit Waldspurger formula [9] the normalised $L$-value $L(1/2,\pi_{\lambda,K}\otimes\chi_{0}\nu)/\Omega_{\lambda}$ is $\ell$-integral.
A key is the following $(\ell,p)$ non-vanishing.
@@
Let $\lambda$ be a self-dual Hecke character over $K$ and $\pi_{\lambda}$ the associated cuspidal automorphic representation of conductor $\mathfrak{N}_{\lambda}$. Let $\ell$ be a prime such that $(\ell,6D_{K})=1$ and ${\mathfrak{p}}\nmid\ell\mathfrak{N}_{\lambda}^{-}$ a prime of $F$. Let $\chi_{0}$ be a finite order anticyclotomic character such that the condition (D) holds. Then for all except finitely many $\nu\in\Xi_{\lambda\chi_{0},{\mathfrak{p}}}^{+}$, we have
@@
since the normalised $L$-values $\pi^{\kappa}L(1,\lambda\chi_{0}\nu)/\Omega_{\infty}^{\Sigma+2\kappa}$ and $\pi^{\kappa}L(1,\lambda(\chi_{0}\nu)^{-1})/\Omega_{\infty}^{\Sigma+2\kappa}$ are algebraic and non-zero for some $\nu\in\Xi_{\lambda\chi_{0},{\mathfrak{p}}}^{+}$. The $\ell$-adic valuation of this ratio of periods is given by the following.
@@
Let $\lambda$ be a self-dual Hecke character over a CM field $K$ of infinity type $\Sigma+\kappa(1-c)$ for $\Sigma$ a CM type of $K$ and $\kappa\in{\mathbb{Z}}_{\geq 0}[\Sigma]$. Let $\ell\nmid 6D_{F}$ be a prime such that $\Sigma$ is an $\ell$-ordinary CM type. Then
The above two results666taking $\chi_{0}=1$ in Theorem 1.11 yield Theorem 1.4 in light of the factorisation (1.3) of $L$-values, and the lower bound for $\ell$-adic valuation of Hecke $L$-values as in  Theorem 5.2.
@@
The above strategy also leads to determination of the $\mu$-invariant of anticyclotomic Katz $p$-adic $L$-function associated to a self-dual Hecke character (cf. [6, Thms. 1.3 &\& 1.5]), giving a different proof of results of [24, 27].
@@
This non-vanishing of Rankin–Selberg $L$-values is based on explicit Waldspurger formula, refined Fourier analysis and equidistribution of special points. For simplicity of notation, we describe the case of $\lambda$ with infinity type $\Sigma$.
@@
for $a]\in{{\rm G}_{n,{\mathfrak{c}}}}:={\mathrm{Gal}}(H_{{\mathfrak{c}}{\mathfrak{p}}^{n}}/K)$ and $H_{{\mathfrak{m}}}$ the ring class field of conductor $\mathfrak{m}$}. By the explicit Waldspurger formula of Cai–Shu–Tian [9], the mod $\ell$ non-vanishing of $L$-values in Theorem 1.11 is equivalent to that of toric periods
@@
where $\nu$ factors through ${\rm G}_{{\mathfrak{c}},n}$. The generality of the work of Cai–Shu–Tian [9] is essential in our study since $D_{K/F}$ divides the conductor of $\pi_{\lambda}$ and $B$ is ramified at primes dividing $D_{K/F}$.
We study $\ell$-indivisibility of the toric periods $P_{f_{\lambda}}(\nu)$ via Fourier analysis and equidistribution of images of special points $x_{{\mathfrak{c}},n}(a)$ in a self-product of the Shimura set $X_{U}$, an idea originally due to Vatsal [57]. The equidistribution is a consequence of Ratner’s seminal ergodicity of unipotent flows  [46]. In light of our refined Fourier analysis on the ring class groups $G_{{\mathfrak{c}},n}$ (see more below), it thus suffices to show that $f_{\lambda}\ \mathrm{mod}\ {\ell}$ is non-Eisenstein on specific components of $X_{U}$. The non-Eisenstein arguments in the work of Vatsal [57] and Cornut–Vatsal [13] do not apply [^7] Note that Cornut–Vatsal [13] is a generalization of Vatsal’s work over totally real fields and essentially excludes the CM case, referred to as exceptional therein. to the CM setting.
@@
with $\rm N$} the reduced norm. For $\mathfrak{p}$} inert in $K$, $f_{\lambda}\ \mathrm{mod}\ {\ell}$ is non-Eisenstein on exactly one of the components $X_{U}^{\pm}$ depending on $\varepsilon(\lambda)$, unlike the work of Vatsal and Cornut–Vatsal where the test vector is non-Eisenstein on all the components. Our non-Eisenstein argument is indirect: $f_{\lambda}\ \mathrm{mod}\ {\ell}$ is non-zero by definition, and consequently non-Eisenstein on $X_{U}$ (see Lemma 4.9 which is specific to the CM setting). So it is non-Eisenstein on at least one of the components $X_{U}^{\tilde{\varepsilon}}\in\{X_{U}^{+},X_{U}^{-}\}$. As the non-vanishing strategy applies on $X_{U}^{\tilde{\varepsilon}}$ and $L(1/2,\pi_{\lambda,K}\otimes\chi_{0}\nu)=0$ for $\nu\in\Xi_{\lambda\chi_{0},{\mathfrak{p}}}^{-}$, we deduce that $\tilde{\varepsilon}$ has the desired parity. Since the test vector $f_{\lambda}$ is independent of $\mathfrak{p}$}, the $\mathfrak{p}$} inert case allows us to treat the $\mathfrak{p}$} split case!
As for the uniform non-vanishing in Theorem 1.11, our decisive new tool: refined Fourier analysis on the ring class groups $G_{{\mathfrak{c}},n}$ via systematic use of the Tunnell–Saito theorem. Indeed the prior Fourier analysis arguments [57, 13, 21, 25] yield the non-vanishing only for a Zariski dense subset of characters if $\deg{\mathfrak{p}}>1$. We partition the domain $\Xi_{\lambda\chi_{0},{\mathfrak{p}}}^{+}$ into subfamilies indexed by certain root of unity and ${\mathbb{Z}}_{p}$-quotients of $\Gamma_{\mathfrak{p}}\simeq{\mathbb{Z}}_{p}^{\deg{{\mathfrak{p}}}}$, and establish uniform non-vanishing over each subfamily. Though there are infinitely many ${\mathbb{Z}}_{p}$-quotients of $\Gamma_{\mathfrak{p}}$ if $\deg{\mathfrak{p}}>1$, we show that they correspond to only finitely many test vectors, leading to the uniform non-vanishing (see §4.5).
@@
For weight two elliptic newforms with square-free conductor, such a comparison is a consequence of Ribet’s level raising (cf. [43, 45]). It may also be approached via ‘$R=T$’ theorems under similar hypothesis (cf. [11, 35, 7]). These methods apply to Hecke eigenforms, and require additional hypotheses over totally real fields. However, our CM setting is neither semistable nor does it involve an eigenform on the Shimura set. It is also excluded by the conjectures of Ichino and Prasanna [45, 31] which pertain to the arithmetic of Petersson norms under the Jacquet–Langlands correspondence.
@@
    an auxiliary degree one prime ${\mathfrak{r}}\nmid{\mathfrak{q}}2\ell{\rm N}_{K/F}({\mathrm{Cond}}(\lambda))$ of $F$ split in $K$ with ${\rm N}_{F/{\mathbb{Q}}}({\mathfrak{r}})\gg 0$ and two characters $\nu_{0},\nu\in\Xi_{\lambda,{\mathfrak{r}}}^{+}$ such that
@@
Since $\varepsilon(\lambda\chi_{0})=\varepsilon(\lambda\chi_{0}^{-1})=+1$, the non-vanishing (1.5) holds by a recent result of the second-named author [20] based on Hida’s strategy (cf. [26]). More precisely, for any prime ${\mathfrak{r}}\nmid 2\ell$ of $F$ split in $K$, [20, Thm. 1.5] for Hecke characters $\lambda\chi_{0}$ and $\lambda\chi_{0}^{-1}$ provides the existence of $\nu_{1},\nu_{2}\in\Xi_{{\mathfrak{r}}}$ such that the non-vanishing (1.5) holds for
@@
where $\sqrt{\nu_{i}}\in\Xi_{{\mathfrak{r}}}$ such that $\sqrt{\nu_{i}}^{2}=\nu_{i}$.
@@
    An $\ell$-integrality of Rankin–Selberg $L$-values: for $\nu\in\Xi_{\lambda\chi_{0},{\mathfrak{r}}}^{+}$ we have
@@
    if ${\mathrm{ord}}_{{\mathfrak{r}}}\mathfrak{N}_{\lambda}<{\mathrm{ord}}_{{\mathfrak{r}}}\mathrm{cond}(\nu_{0,{\mathfrak{r}}})\leq{\mathrm{ord}}_{{\mathfrak{r}}}\mathrm{cond}(\nu_{{\mathfrak{r}}})$, where $\nu_{{\mathfrak{r}}}:=\nu|_{K_{\mathfrak{r}}^{\times}}$ and $\mathrm{cond}(\nu_{{\mathfrak{r}}})$ is the maximal integral ideal ${\mathfrak{r}}^{k}$ of $O_{\mathfrak{r}}$ such that $\nu_{\mathfrak{r}}((O_{\mathfrak{r}}+{\mathfrak{r}}^{k}O_{K_{\mathfrak{r}}})^{\times})=1$.
@@
    The mod $\ell$ non-vanishing of Rankin–Selberg $L$-values: Theorem 1.11 for the base character $\lambda\nu_{0}$, prime $\mathfrak{r}$} and $\chi_{0}\in\Xi_{\lambda,{\mathfrak{q}}}^{+}$.
@@
    A lower bound for $\ell$-divisibility of Hecke $L$-values: for $\nu\in\Xi_{\lambda\chi_{0},{\mathfrak{r}}}^{+}$ we have

### v2

@@ 1.4. Strategy
The non-vanishing problems (Q1) and (Q2) concerns the arithmetic of a self-dual Hecke character or that of the unitary group $U(1)$ over $K$. Via theta correspondence, we recast them as problems on a totally definite unitary group $U(2)$ over $K$. The latter turn out to be amenable to automorphic and ergodic tools.
@@ 1.4.1. Shimura set
For a CM field $K$ and $F$ the maximal totally real subfield, we have a canonically associated totally definite quaternion algebra $B$ over $F$ characterised by
@@
for any place $v$ of $F$. Note that $K$ embeds into $B$ as an $F$-algebra, and we fix such an embedding.
@@ 1.4.2. Ancillary results
Our mod $\ell$ non-vanishing of Hecke $L$-values is based on Theorems 1.9 and 1.10 below, which concern the arithmetic of the Shimura set.
For a self-dual Hecke character $\lambda$ over $K$ of infinity type $\Sigma+\kappa(1-c)$, let $\phi_{\lambda}$ be the associated ${\mathrm{GL}}_{2}(F)$-theta series of weight $k:=2\Sigma+2\kappa$ and level $U_{0}(\mathfrak{N}_{\lambda})$ for $\mathfrak{N}_{\lambda}:=D_{K/F}{\rm N}_{K/F}({\mathrm{Cond}}(\lambda))$. Write $\mathfrak{N}_{\lambda}=\mathfrak{N}_{\lambda}^{+}\mathfrak{N}_{\lambda}^{-}$ with $\mathfrak{N}_{\lambda}^{+}$ (resp. $\mathfrak{N}_{\lambda}^{-}$) only divisible by split (resp. non-split) primes of $F$ in $K$. Let $\pi_{\lambda}$ be the Jacquet–Langlands transfer of the automorphic representation generated by $\phi_{\lambda}$ to $B^{\times}$. Let $\ell$ be a prime such that $(\ell,6{\mathfrak{p}}D_{K})=1$. Let
@@
be the CM form as in subsection 3.4. It is a test vector in the sense of Gross and Prasad [17], which is $\ell$-primitive and $K_{v}^{\times}$-invariant for all places $v|\mathfrak{N}_{\lambda}^{-}$. We emphasise that $f_{\lambda}$ is not a newform [^4] In this setting the finite part of the discriminant of $B$ divides $D_{K/F}$. Consequently, at places dividing $D_{K/F}$, the newform is not a test vector under an optimal embedding $K\hookrightarrow B$..
@@
The Rankin–Selberg $L$-values $L(1/2,\pi_{\lambda,K}\otimes\chi_{0}\nu)$ maybe studied via the Waldspurger formula. While employing it, the period
@@
appears naturally, where $(\phi_{\lambda},\phi_{\lambda})_{U_{0}(\mathfrak{N}_{\lambda})}$ and $\langle{f_{\lambda},f_{\lambda}}\rangle_{\widehat{R}^{\times}}$ are Petersson norms. In view of the explicit Waldspurger formula [8] the normalised $L$-value $L(1/2,\pi_{\lambda,K}\otimes\chi_{0}\nu)/\Omega_{\lambda}$ is $\ell$-integral.
A key is the following $(\ell,p)$ non-vanishing (cf. Theorems 4.15 and 4.19).
@@
Let $\lambda$ be a self-dual Hecke character over $K$ and $\pi_{\lambda}$ the associated cuspidal automorphic representation of conductor $\mathfrak{N}_{\lambda}$. Let $\ell$ be a prime such that $(\ell,6D_{K})=1$ and ${\mathfrak{p}}\nmid\ell\mathfrak{N}_{\lambda}^{-}$ a degree one prime of $F$. Let $\chi_{0}$ be a finite order anticyclotomic character such that the condition (D) holds. Then for all except finitely many $\nu\in\Xi_{\lambda\chi_{0},{\mathfrak{p}}}^{+}$, we have
@@
since the normalised $L$-values $\pi^{\kappa}L(1,\lambda\chi_{0}\nu)/\Omega_{\infty}^{\Sigma+2\kappa}$ and $\pi^{\kappa}L(1,\lambda(\chi_{0}\nu)^{-1})/\Omega_{\infty}^{\Sigma+2\kappa}$ are algebraic and non-zero for some $\nu\in\Xi_{\lambda\chi_{0},{\mathfrak{p}}}^{+}$. The $\ell$-adic valuation of this ratio of periods is given by the following (cf. Theorem 5.6).
@@
Let $\lambda$ be a self-dual Hecke character over a CM field $K$ of infinity type $\Sigma+\kappa(1-c)$ for $\Sigma$ a CM type of $K$ and $\kappa\in{\mathbb{Z}}_{\geq 0}[\Sigma]$. Let $\ell\nmid 6D_{F}$ be a prime such that $\Sigma$ is an $\ell$-adic CM type. Then
The above two results555taking $\chi_{0}=1$ in Theorem 1.9 yield Theorem 1.4 in light of the factorisation (1.3) of $L$-values, and the lower bound for $\ell$-adic valuation of Hecke $L$-values as in  Theorem 5.2.
@@
The above strategy also leads to determination of the $\mu$-invariant of anticyclotomic Katz $p$-adic $L$-function associated to a self-dual Hecke character (cf. [5, Thms. 1.3 &\& 1.5]), giving a different proof of results of [23, 26].
@@
This mod $\ell$ non-vanishing of the Rankin–Selberg $L$-values $L^{{\mathrm{alg}}}(1/2,\pi_{\lambda,K}\otimes\chi_{0}\nu):=L(1/2,\pi_{\lambda,K}\otimes\chi_{0}\nu)/\Omega_{\lambda}$ is based on explicit Waldspurger formula, Fourier analysis on ring class groups in conjunction with automorphic representation theory and equidistribution of special points. For simplicity of notation, we describe the case of $\lambda$ with infinity type $\Sigma$.
@@
for $a]\in{{\rm G}_{n,{\mathfrak{c}}}}:={\mathrm{Gal}}(H_{{\mathfrak{c}}{\mathfrak{p}}^{n}}/K)$ and $H_{{\mathfrak{m}}}$ the ring class field of conductor $\mathfrak{m}$}. Based on explicit Waldspurger formula of Cai–Shu–Tian (cf. Theorem 3.14) and test vector theory (cf. §3.2), the mod $\ell$ non-vanishing of $L^{{\mathrm{alg}}}(1/2,\pi_{\lambda,K}\otimes\chi_{0}\nu)$ is equivalent to that of toric periods
@@
where $\nu$ factors through ${\rm G}_{{\mathfrak{c}},n}$ (cf. §4.3) . The generality of the work of Cai–Shu–Tian [8] is essential in our study since $D_{K/F}$ divides the conductor of $\pi_{\lambda}$ and $B$ is ramified at primes dividing $D_{K/F}$.
We study $\ell$-indivisibility of the toric periods $P_{f_{\lambda}}(\nu)$ via automorphic representation theory, Fourier analysis on the ring class groups ${\rm G}_{{\mathfrak{c}},n}$ and equidistribution of images of special points $x_{{\mathfrak{c}},n}(a)$ in a self-product of the Shimura set $X_{U}$. The equidistribution is a consequence of Ratner’s seminal ergodicity of unipotent flows  [44]. The relevance of equidisitribution to mod $\ell$ non-vanishing is an insight of Vatsal [55]. His strategy was generalised to totally real fields by Cornut–Vatsal [12]. However, their results essentially excludes the CM case, referred to as exceptional therein. To approach the CM case, we refine the strategy by considering toric periods over Galois orbits and by employing automorphic representation theory peculiar to the CM case (see more below). It then suffices to show that $f_{\lambda}\ \mathrm{mod}\ {\ell}$ is non-Eisenstein on specific components of $X_{U}$, which encode the $\varepsilon$-factor $\varepsilon(\lambda)$. The non-Eisenstein arguments in the work of Vatsal [55] and Cornut–Vatsal [12] do not apply to our CM setting.
@@
with $\rm N$} the reduced norm. For $\mathfrak{p}$} inert in $K$, $f_{\lambda}\ \mathrm{mod}\ {\ell}$ is non-Eisenstein on exactly one of the components $X_{U}^{\pm}$ depending on $\varepsilon(\lambda)$, unlike the work of Vatsal and Cornut–Vatsal where the test vector is non-Eisenstein on all the components. Our non-Eisenstein argument is indirect: $f_{\lambda}\ \mathrm{mod}\ {\ell}$ is non-zero by definition, and consequently non-Eisenstein on $X_{U}$ (see Lemma 4.10 which is specific to the CM setting). So it is non-Eisenstein on at least one of the components $X_{U}^{\tilde{\varepsilon}}\in\{X_{U}^{+},X_{U}^{-}\}$. As the non-vanishing strategy applies on $X_{U}^{\tilde{\varepsilon}}$ and $L(1/2,\pi_{\lambda,K}\otimes\chi_{0}\nu)=0$ for $\nu\in\Xi_{\lambda\chi_{0},{\mathfrak{p}}}^{-}$, we deduce that $\tilde{\varepsilon}$ has the desired parity. Since the test vector $f_{\lambda}$ is independent of $\mathfrak{p}$}, the $\mathfrak{p}$} inert case allows us to treat [^6] See §4.4.2. the $\mathfrak{p}$} split case!
We now elaborate uniform non-vanishing of the toric periods $P_{f_{\lambda}}(\nu)$ (cf. §4.4). As $\nu$ varies in a Galois orbit, the average of toric periods $P_{f_{\lambda}}(\nu)$ leads to a partial toric period over a thin subgroup ${\rm G}_{{\mathfrak{c}},n}^{\prime}\subset{\rm G}_{{\mathfrak{c}},n}$ for $n:={\mathrm{ord}}_{\mathfrak{p}}\mathrm{cond}(\nu)$, whose cardinality is independent of $n$ since $\deg{\mathfrak{p}}=1$ (see (4.10) and Remark 4.16). We analyse the period over ${\rm G}_{{\mathfrak{c}},n}^{\prime}$ by partitioning ${\rm G}_{{\mathfrak{c}},n}^{\prime}$ into a locally defined subgroup ${\rm H}_{{\mathfrak{c}},n}$ and $D_{1}\subset\widehat{K}^{\times}$, the latter representing ${\rm G}_{{\mathfrak{c}},n}^{\prime}/{\rm H}_{{\mathfrak{c}},n}$. The corresponding periods are studied via local representation theory and ergodic theory respectively, as outlined below.
We first introduce a partition of $\Xi_{\mathfrak{p}}$ into finitely many subfamilies $\mathcal{S}$} (see the discussion above (4.24)) and a Hecke-modification $f_{{\mathcal{S}}}$ of the test vector $f_{\lambda}$ as in (4.24) such that the ${\rm H}_{{\mathfrak{c}},n}$-action has the following uniform eigen-behaviour:
  • $\bullet$
    $f_{{\mathcal{S}}}(x_{{\mathfrak{c}},n}(\cdot))$ is $(\chi_{0}\nu)^{-1}$-eigen under the ${\rm H}_{{\mathfrak{c}},n}$-action for each $n$ with $\nu\in{\mathcal{S}}$ and $n={\mathrm{ord}}_{\mathfrak{p}}\mathrm{cond}(\nu)$.
By the Tunnell–Saito theorem and its mod $\ell$ variant, $f_{{\mathcal{S}}}$ is still a test vector for the subfamily $\mathcal{S}$} and has the same $\ell$-divisibility as $f_{\lambda}$ (cf. Lemmas 3.12 and 4.8). In view of the eigen-property, the toric period of $f_{\lambda}$ equals that of $f_{\mathcal{S}}$, and so in our strategy the former maybe replaced by the latter. A key feature of $f_{{\mathcal{S}}}$: the associated toric period over ${\rm G}_{{\mathfrak{c}},n}^{\prime}$ factors through ${\rm G}_{{\mathfrak{c}},n}^{\prime}/{\rm H}_{{\mathfrak{c}},n}$ . Hence, its non-vanishing can be studied via the equidistribution of skewed-diagonal images of special points on the $D_{1}-$fold self-product of the Shimura set $X_{U}$, leading to uniform non-vanishing of the toric periods over the subfamily $\mathcal{S}$} (cf. Lemma 4.2 and Proposition 4.5). Since there are finitely many subfamilies, Theorem 1.9 follows.
@@
For weight two elliptic newforms with square-free conductor, such a comparison of periods is a consequence of Ribet’s seminal level raising (cf. [41, 43]). For elliptic newforms with weights in the Fontaine–Laffaille range and square-free conductors, it may also be approached via ‘$R=T$’ theorems (cf. [10, 33, 6]). These methods apply to Hecke eigenforms, and require additional hypotheses over totally real fields. However, our CM setting is neither semistable nor does it involve an eigenform on the Shimura set, and the weights are not necessarily in the Fontaine–Laffaille range [^7] As far as we know, our result is the first result in the non Fontaine–Lafaille range.. It is also excluded by the conjectures of Ichino and Prasanna [43, 30] which pertain to the arithmetic of Petersson norms under the Jacquet–Langlands correspondence.
@@
    an auxiliary degree one prime ${\mathfrak{q}}^{\prime}\nmid{\mathfrak{q}}2\ell{\rm N}_{K/F}({\mathrm{Cond}}(\lambda))$ of $F$ split in $K$ with ${\rm N}_{F/{\mathbb{Q}}}({\mathfrak{q}}^{\prime})\gg 0$ and two characters $\nu_{0},\nu\in\Xi_{\lambda,{\mathfrak{q}}^{\prime}}^{+}$ such that
@@
Since $\varepsilon(\lambda\chi_{0})=\varepsilon(\lambda\chi_{0}^{-1})=+1$, the non-vanishing (1.4) holds by a recent result of the second-named author [19] based on Hida’s strategy (cf. [25]). More precisely, for any prime ${\mathfrak{q}}^{\prime}\nmid 2\ell$ of $F$ split in $K$, [19, Thm. 1.5] for Hecke characters $\lambda\chi_{0}$ and $\lambda\chi_{0}^{-1}$ provides the existence of $\nu_{1},\nu_{2}\in\Xi_{{\mathfrak{q}}^{\prime}}$ such that the non-vanishing (1.4) holds for
@@
where $\sqrt{\nu_{i}}\in\Xi_{{\mathfrak{q}}^{\prime}}$ such that $\sqrt{\nu_{i}}^{2}=\nu_{i}$.
@@
    An $\ell$-integrality of Rankin–Selberg $L$-values: for $\nu\in\Xi_{\lambda\chi_{0},{\mathfrak{q}}^{\prime}}^{+}$ we have
@@
    if ${\mathrm{ord}}_{{\mathfrak{q}}^{\prime}}\mathfrak{N}_{\lambda}<{\mathrm{ord}}_{{\mathfrak{q}}^{\prime}}\mathrm{cond}(\nu_{0,{\mathfrak{q}}^{\prime}})\leq{\mathrm{ord}}_{{\mathfrak{q}}^{\prime}}\mathrm{cond}(\nu_{{\mathfrak{q}}^{\prime}})$, where $\nu_{{\mathfrak{q}}^{\prime}}:=\nu|_{K_{{\mathfrak{q}}^{\prime}}^{\times}}$ and $\mathrm{cond}(\nu_{{\mathfrak{q}}^{\prime}})$ is the maximal integral ideal $({\mathfrak{q}}^{\prime})^{k}$ of $O_{{\mathfrak{q}}^{\prime}}$ such that $\nu_{{\mathfrak{q}}^{\prime}}((O_{{\mathfrak{q}}^{\prime}}+{{\mathfrak{q}}^{\prime}}^{k}O_{K_{{\mathfrak{q}}^{\prime}}})^{\times})=1$.
@@
    The mod $\ell$ non-vanishing of Rankin–Selberg $L$-values: Theorem 1.9 for the base character $\lambda\nu_{0}$, prime $\mathfrak{q}$$^{\prime}$ and $\chi_{0}\in\Xi_{\lambda,{\mathfrak{q}}}^{+}$.
@@
    A lower bound for $\ell$-divisibility of Hecke $L$-values: for $\nu\in\Xi_{\lambda\chi_{0},{\mathfrak{q}}^{\prime}}^{+}$ we have